OpenStudy (anonymous):

y varies directly as the square of x. When x = 1, y = 3. What is the value of y when x = 3?

OpenStudy (anonymous):

y = kx^2 k = 3 y = (3)*9 = 27

OpenStudy (anonymous):

y = kx^2 find k

OpenStudy (anonymous):

27

OpenStudy (anonymous):

thankk uuu :D

OpenStudy (amistre64):



OpenStudy (anonymous):

amistre, y varies with x squared, not just with x.

OpenStudy (amistre64):

.... lol, well that changes things a bit ;)

OpenStudy (anonymous):

indeed it does :)

OpenStudy (amistre64):